Find an antiderivative F(x) of f(x) = 2 − 3e^x so that F(1) = 4. show work
Answer #1

TO tons f(x) = 2-3e* Antiderivative of fre) is. F(x)= Speech = 5(2-384)dx = 2x=3&* te, e being integrating constant Now F (i)